OpenStudy (anonymous):

What is the solution to sqrt(1-4x)=3?

OpenStudy (anonymous):

\[\sqrt{1-4x}=3\] first you got to get rid of that square root by squaring everything 1-4x=9 minus 1 -4x=8 divide by -4 x=-2 substitute to see if it works

OpenStudy (anonymous):

yes it worked,thank you so much :)

OpenStudy (haseeb96):

\[\sqrt{1-4x}=3\] taking square on both sides now it becomes 1-4x=9 1-9=4x -8\4=x x=-2 now for verification put x=-2 in the given equation \[\sqrt{1-4(-2)}\] \[\sqrt{1+8}=3\] \[\sqrt{9}=3\] 3=3 hence shown